Sora Industries has 6868 million outstanding shares, $120 million in debt, $49 million in cash, and the following projected free cash flow for the next four years :

Year 0 1 2 3 4
Earning & FCF Forecast ($millions)
1 Sales 433 468 516 547 574.3
2    Growth vs. Prior Year 8.1% 10.3% 6.0% 5.0%
3 Cost of Goods Sold (313.6) (345.7) (366.5) (384.8)
4 Gross Profit 154.4 170.3 180.5 189.5
5 Selling, General & Admin. (93.6) (103.2) (109.4) (114.9)
6 Depreciation (7.0) (7.5) (9.0) (9.5)
7 EBIT 53.8 59.6 62.1 65.2
8 Less: Income tax at 40% (21.5) (23.8) (24.8) (26.1)
9 Plus: Depreciation 7 7.5 9 9.5
10 Less: Capital Expenditures (7.7) (10.0) (9.9) (10.4)
11 Less: Increases in NWC (6.3) (8.6) (5.6) (4.9)
12 Free Cash Flow 25.3 24.6 30.8 33.3

a. Suppose Sora's revenue and free cash flow are expected to grow at a 4.4% rate beyond year 4. If Sora's weighted average cost of capital is 11.0% , what is the value of Sora's stock based on this information?

b. Sora's cost of goods sold was assumed to be 67% of sales. If its cost of goods sold is actually 70% of sales, how would the estimate of the stock's value change?

c. Let's return to the assumptions of part (a ) and suppose Sora can maintain its cost of goods sold at 67% of sales. However, now suppose Sora reduces its selling, general, and administrative expenses from 20% of sales to 16% of sales. What stock price would you estimate now? (Assume no other expenses, except taxes, are affected.)

d. Sora's net working capital needs were estimated to be 18% of sales (which is their current level in year 0). If Sora can reduce this requirement to 12% of sales starting in year 1, but all other assumptions remain as in part (a ), what stock price do you estimate for Sora? (Hint: This change will have the largest impact on Sora's free cash flow in year 1.)
Answer #1

Case (a)

Step 1: Calculating Terminal Value (TV)

TV = jFQxSDtnGn1oF5eKMPgl0ANpdgeO0kIojVYHjEYl

FCFt = last year's projected Free Cash Flow = $ 33.3 million or 33.3

g = expected growth rate = 4.4% or 0.044

r = cost of capital = 11% or 0.11

TV = AQiubmWG+5xx0jlnLnPqMuffusAF+EIIADs=

TV = cvBrfhjQ6EQulNLT+bET59e8OqCWe7pIc2HCbJVR

TV = 526.7455

Step 2: Calculating Present Value of Terminal Value (PVTV)

PVTV =NAirnOPCIFjAfSCt+XvW8deDMWYyOgKwWANeCYc0

r = Cost of Capital = 11% or 0.11%

t = time period = 4

PVTV =Wg4aiCkecZOzaN389TGuyTfaZqFHTwn87ld6PbbN = 346.9835

Step 3: Calculating Present Value of Projected Free Cash Flows(PVFCF)

PVFCF = Zw9T6JqJffsSbQhL4QochMhAAA7

Ci = Projected Free Cash Flows

r = Cost of Capital = 11% or 0.11

PV = pIADSx8B9BGvzyScvGoiHFNhFLEGP4cCA2i+16Rd + 16G8MtEayxsOhqLSgCBPlC4cV0GNEgIAOw== + VN92YtrFkIoMTK65RGUgEIZtKYuwrQs3cNXS3eBB + T8lAEBmSkSTG5KytFGCAA7 = 87.22

Step 4: Calculating Net Present Value of Firm(PVF)

PVF = PVTV + PVFCF = 346.9835 +87.22 =$ 434.2035 million

Step 5: Subtracting Value of Debt from Net Present Value of Firm

Value of Debt = $ 120 million

Net Present Value of Firm = $ 434.2035

434.2035 - 120 = $ 314.2035 million = Value of Equity

Step 5: Calculating Value of Stock

Value of Stock = Value of Equity / Number of Equity Shares Outstanding

Value of Stock = 314.2035 / 6868 = $ 0.045749

Case (b)

New FCF statement after adjusting COGS being 70% of Sales

0

1

2

3

4

1

Sales

433.00

468.00

516.00

547.00

574.30

3

Cost of Goods Sold

-327.60

-361.20

-382.90

-402.01

4

Gross Profit

140.40

154.80

164.10

172.29

5

Selling, General & Admin.

   -93.60

-103.20

-109.40

-114.90

6

Depreciation

     -7.00

     -7.50

     -9.00

     -9.50

7

EBIT

    39.80

    44.10

    45.70

    47.89

8

Less: Income tax at 40%

   -15.92

   -17.64

   -18.28

   -19.16

9

Plus: Depreciation

      7.00

      7.50

      9.00

      9.50

10

Less: Capital Expenditures

     -7.70

   -10.00

     -9.90

   -10.40

11

Less: Increases in NWC

     -6.30

     -8.60

     -5.60

     -4.90

12

Free Cash Flow

    16.88

    15.36

    20.92

    22.93

Using case (a) approach and using new FCF,

Terminal Value = 362.7109

Present Value of Terminal Value = 238.9289

Present Value of Projected Free Cash Flows = 58.08

Net Present Value of Firm = 297.01

Subtracting Value of Debt from Net Present Value of Firm = 177.01

Value of Stock = $ 0.025772

Change in Stock Value in comparison to case (a) = 0.025772 - 0.045749 = -0.019976 or -43.66%

Case (c)

New FCF statement after adjusting selling, general, and administrative expenses being 16% of sales

0

1

2

3

4

1

Sales

433.00

468.00

516.00

547.00

574.30

3

Cost of Goods Sold

-313.6

-345.7

-366.5

-384.8

4

Gross Profit

154.4

170.3

180.5

189.5

5

Selling, General & Admin.

-74.88

-82.56

-87.52

-91.888

6

Depreciation

-7

-7.5

-9

-9.5

7

EBIT

72.52

80.24

83.98

88.112

8

Less: Income tax at 40%

-29.008

-32.096

-33.592

-35.245

9

Plus: Depreciation

7

7.5

9

9.5

10

Less: Capital Expenditures

-7.7

-10

-9.9

-10.4

11

Less: Increases in NWC

-6.3

-8.6

-5.6

-4.9

12

Free Cash Flow

36.512

37.044

43.888

47.0672

Using case (a) approach and using new FCF,

Terminal Value = 744.5175

Present Value of Terminal Value = 490.4368

Present Value of Projected Free Cash Flows = 126.05

Net Present Value of Firm = 616.49

Subtracting Value of Debt from Net Present Value of Firm = 496.49

Value of Stock = $ 0.072290

Change in Stock Value in comparison to case (a) = 0.072290- 0.045749 = 0.026541 or 58.01%